OpenStudy (anonymous):

What is happening if energy input remains constant and voltage remains the same in a circuit, but the current decreases? A. The resistance has increased. B. The volts have decreased. C. The power has increased.

OpenStudy (anonymous):

im checking my textbook right now

OpenStudy (anonymous):

i want to say it is A or C

OpenStudy (anonymous):

i think its A

OpenStudy (anonymous):

Use the various expressions for electrical power in terms of voltage \(V\), current \(I\) and resistance \(R\) \[P=VI=I^2R=\frac{V^2}{R}\] Power is constant, voltage is constant, current is decreasing, there's only one thing left that can change.

OpenStudy (anonymous):

It sorta tells you in the question that 2 of the answers are wrong. "Energy input is the same" "voltage remains the same"

OpenStudy (anonymous):

A because it is resistance has increased

OpenStudy (anonymous):

can i get a badge plz :)

OpenStudy (anonymous):

It has to be A... although I think there's something wrong with the question. Hang on a second.

OpenStudy (anonymous):

thxx

OpenStudy (anonymous):

thank you both

OpenStudy (anonymous):

\(P = I^2R\) remains constant \(V=IR\) remains constant If \(I\) decreases by half \(I\to\frac12 I\) then in order to conserve power, \(R \to 4R\) so \(P = (\frac12 I)^2\times4 R = I^2R\) and in order to conserve voltage \(R \to 2R\) so \(V=\frac12 I\times2R = IR\) Resistance cannot both double and quadruple. The question describes an impossible circuit.